OpenStudy (anonymous):

Question 2.2. The temperature of an oven went from 330° to 396°. What was the percent increase in temperature? (Points : 5) 2% 16.6% 20% 66%

OpenStudy (perl):

$$ \Large \rm \left|\frac{new - original }{original }\right| \times 100 = Percent ~ Change $$

OpenStudy (anonymous):

can i give the direct answer

OpenStudy (anonymous):

or no

OpenStudy (perl):

better guide the student

OpenStudy (anonymous):

66

OpenStudy (perl):

sid thats incorrect

OpenStudy (perl):

$$ \Large \rm \left|\frac{new - original }{original }\right| \times 100 = Percent ~ Change \\~\\~\\ \Large \rm \left|\frac{396 - 330 }{330 }\right| \times 100 $$

OpenStudy (anonymous):

this is the worst subject i suck badly so i dont get it could you help me to understand it

OpenStudy (perl):

sure just do the division first

OpenStudy (anonymous):

first multiply each percentage by 396 and then the answer you get from multiplying subtract it

OpenStudy (perl):

$$ \Large \rm \frac{new - original }{original } \times 100 = Percent ~ Change \\~\\~\\ \Large \rm \frac{396 - 330 }{330 }\times 100 \\~\\~\\ \Large \rm \frac{66 }{330 }\times 100 $$

OpenStudy (anonymous):

for example

OpenStudy (anonymous):

20

OpenStudy (perl):

@unkownbeast911 to make your method work you have to multiply each percentage by 330, not 396

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Perl can you help me with some not telling me the answer but helping me understand it

OpenStudy (perl):

then subtract the result from 396, and see if it matches 330

OpenStudy (perl):

ok sid

OpenStudy (anonymous):

want me to open a new question or message ?

OpenStudy (perl):

lets open a new question

OpenStudy (anonymous):

ok

OpenStudy (perl):

in case i get disconnected, other people can help